# OntoPortal-Astro, a Semantic Artefact Catalogue for Astronomy

### v1

@@ Abstract
The astronomy communities are widely recognised as mature communities for their open science practices. However, while their data ecosystems are rather advanced and permit efficient data interoperability, there are still gaps between these ecosystems. Semantic artefacts (e.g., ontologies, thesauri, vocabularies or metadata schemas) are a means to bridge that gap as they allow to semantically described the data and map the underlying concepts. The increasing use of semantic artefacts in astronomy presents challenges in description, selection, evaluation, trust, and mappings. The landscape remains fragmented, with semantic artefacts scattered across various registries in diverse formats and structures — not yet fully developed or encoded with rich semantic web standards like OWL or SKOS — and often with overlapping scopes. Enhancing data semantic interoperability requires common platforms to catalog, align, and facilitate the sharing of FAIR semantic artefacts. In the frame of the FAIR-IMPACT project, we prototyped a semantic artefact catalogue for astronomy, heliophysics and planetary sciences. This exercise resulted in improved vocabulary and ontology management in the communities, and is now paving the way for better interdisciplinary data discovery and reuse. This article presents current practices in our discipline, reviews candidate SAs for such a catalogue, presents driving use cases and the perspective of a real production service for the astronomy community based on the OntoPortal technology, that will be called OntoPortal-Astro.
@@ 1 Introduction
The set of resources of the so-called “astronomy virtual observatory” is a web of related data products, catalogues, services and tools allowing researchers to access and process data for their research. This is implemented thanks to the International Virtual Observation Alliance111 https://ivoa.net (IVOA) framework (Arviset et al., 2018), which proposes a registry (Demleitner et al., 2019) gathering endpoints for services, datasets and catalogues, and is used by tools to discover and provide access to resources. This astronomy framework is an example of an open linked-data ecosystem.
The linked-data architecture has been defined at the inception times of the world-wide-web (Berners-Lee and Hendler, 2001). The original framework — called RDF: Resource Description Framework — (Schreiber and Raimond, 2014) developed and published by the W3C proposes a description of resources and their relations, based on triples: a subject, a predicate and an object. The triples can be seen as sentences, the predicate being the verb. The set of triples creates an oriented and tagged graph of nodes. Nodes are resources (referred to as URI: Unified Resource Identifier). Predicates (or properties) designate the links connecting the ressources. The classes, their instances and the properties are usually defined in standard vocabularies or ontologies capturing a certain level of semantics that a machine will then rely to query and reason over the data. In the present case, this graph of nodes and relations is capturing knowledge about scientific data, hence we call it a “knowledge graph”. Keeping this knowledge graph of web resources usable and interoperable while it grows requires semantic and technical coordination and synchronisation, so that terms are globally understood by the tools and actors providing the content and enabling the access. This semantic and technical coordination is achieved by methods, technologies and tools coming from a domain called, the Semantic Web for which semantics artefacts, are at the core (Corcho et al., 2024). A semantic artefact (hereafter referred to as SA) — a broader term to include ontologies, terminologies, taxonomies, thesauri, vocabularies, metadata schemas — can be a simple list of terms (properties, classes of resources), or a more complex ontology allowing inference and logics between terms. We know that a unique absolute ontology is utopia. Each SAs encode contextual knowledge, which need to be understandable, shareable and reusable by humans and machines.
@@
The astronomy community is mainly composed of three semantics sub-communities: celestial astronomy (objects are referenced to with their fixed sky coordinates, e.g., stars, galaxies, etc); heliophysics (the study of the Sun, and the plasma environments throughout the Solar System); planetary sciences (the study of the Solar System objects, e.g., planets, comets, asteroids, etc). Each of these sub-communities have developed interoperability frameworks and semantic ecosystems, which are rather siloed up to now, compromising the semantic and technical interoperability of those ecosystems.
@@
The main astronomy sub-community is the celestial astronomy community. It is organised around the IVOA, which is interoperability driven. It defines schemas, protocols and vocabularies, which are respectively managed by the Data Models, Data Access Layers, and Semantics working groups. In addition, a registry (managed by the IVOA Registry working group) provides a catalogue of IVOA compliant resources. Hence, this enables a stable and well-established ecosystem adopted by the vast majority of astronomy archives, data centres and tool developers. As a result, scientists are using applications, such as Aladin (Bonnarel et al., 2000) or TOPCAT (Taylor, 2005, 2011), which are fully relying on IVOA tooling without seeing any detail of the underlying interoperability layers.
@@
The second sub-community is the heliophysics community. It is organised around the International Heliophysics Data Environment Alliance [^2] https://ihdea.net (IHDEA, Masson et al., 2024), which is distribution driven. The Space Physics Archive Search and Extract (SPASE, Roberts et al., 2018) is the main heliophysics information model. It defines a rather complete and evolutive way to describe space physics observational datasets, model runs, instruments or repositories. It is used by several major heliophysics data centres (NASA/SPDF, CNES/CDPP, IUGONET, see e.g., Abe et al., 2014; Clarke, 2016; Roberts et al., 2018) for their internal data model. Other metadata guidelines are available to this community (such as the ISTP metadata for data stored in CDF (Whipple and Lancaster, 1995; Candey, 1998), or the SOLARNET metadata for data stored in FITS (Hanisch et al., 2001; Haugan and Fredvik, 2021)). Data access protocols are also standardised, with, e.g., the HAPI protocol (Weigel et al., 2021). The IHDEA ecosystem is not as interoperable as the IVOA one. While some tools are providing a very simple access to heliophysics datasets (such as the AMDA [^3] https://amda.irap.omp.eu (Génot et al., 2021) or the CDAWeb [^4] https://cdaweb.gsfc.nasa.gov web interfaces), using IHDEA services or resources often require a deep understanding of interoperability interfaces and data formats. However ongoing developments are aiming at building tighter links with IVOA and improve open science in heliophysics.
@@
The third sub-community is the planetary science community. It is mainly organised around the International Planetary Data Alliance [^5] https://planetarydata.org (IPDA), which is archive driven. The primary role of the IPDA is to develop and maintain the NASA Planetary Data System (PDS) (Hughes et al., 2017), which is an advanced information model for planetary exploration science archives. The PDS is used by all major space agency archive with planetary exploration missions: NASA/PDS, ESA/PSA, JAXA, ISRO, etc. Moreover, planetary scientists studying the planetary surfaces are using the OGC (Open Geospatial Consortium) ecosystem, developed by the geoscience community (Earth observation). Technical interoperability between OGC and IVOA has been studied for some years (Hare et al., 2018; Marmo et al., 2018; Minin et al., 2019; Erard and the VESPA team, 2024). However, semantic interoperability still needs to be implemented.
@@
The scope of this study covers the astronomy community along with its three sub-domains (celestial astronomy, solar system sciences and heliophysics). Recent publications (e.g., Hurlburt and Timmons, 2022; Berriman, 2024; Cecconi, 2024; Landais et al., 2024; Servillat et al., 2024; Masson et al., 2024) show that each of these communities are claiming that they are producing a FAIR (Findable, Accessible, Interoperable and Reusable, Wilkinson et al., 2016) ecosystem. The evaluation of the FAIRness of digital resources is discussed and studied by many groups, and is out of the scope of this paper. However, the fact that each sub-community is still rather siloed is a hint that things could be improved. In 2018, the IVOA Solar System interest group started to bring heliophysics and planetary sciences needs in the IVOA ecosystem. The uptake of IVOA protocols in heliophysics and planetary sciences is growing, but at lot remains to be done, in particular on the semantic interoperability side, so that it is more widely adopted. This means that we need to harmonise our SAs, so that communities can exchange data and transform these data into knowledge — semantically described, interoperable, actionable, and open. Moreover, the adoption of SAs across communities will improve the FAIRness of data products, since the metadata would be described with the same terms, or with terms that are related one to the other. The increasing use of SAs in astronomy presents challenges including their description, selection, evaluation, trustworthiness and mapping. The landscape remains fragmented, with SAs scattered across various registries in diverse formats and structures — not yet fully developed or encoded with rich semantic web standards based on RDF like OWL or SKOS — and often with overlapping scopes. There is thus need for a common place to catalogue, manage, assess, select and align FAIR SAs from several communities (Gray, 2015; Ritschel et al., 2016; Rovetto, 2023).
In this paper, we present the results of a prototyping exercise. Thanks to the FAIR-IMPACT project, we have implemented an instance of the OntoPortal technology (Jonquet et al., 2023) to deploy a prototype SA catalogue for astronomy. Hereafter, we present current practices in our discipline, and review candidate SAs for such a cataloguing exercise. We report the use cases that have been driving this development. We present the first results, discuss the impact and outcomes, and layout a roadmap for the deployment of a real production service for the astronomy community based on the OntoPortal technology, that will be called OntoPortal-Astro. Throughout the paper, the astronomy OntoPortal instance and the OntoPortal-Astro instances refer respectively to the prototype instance development within the FAIR-IMPACT project, and the upcoming official server.
@@ 2 Current semantic artefact management in astronomy
The celestial astronomy community data alliance is the IVOA, which maintains an operational interoperability framework used by data repositories and science application platforms throughout the world. In this community, SAs are composed of vocabularies and schemas. The Semantics Working Group of the IVOA is managing the vocabularies used in the IVOA standards, following the Vocabularies in the VO specification (Demleitner et al., 2023). The vocabularies are available from a dedicated web page [^6] https://ivoa.net/rdf/, and are accessible using IVOA and RDF tooling. The Data Model working group of the IVOA is managing the schemas, following the Virtual Observatory Data Model Language specification (Lemson et al., 2018), but they are not available through the RDF tooling.
@@
In the planetary science community, two main frameworks co-exist with different scopes: the IPDA, which proposes an advanced data archiving metadata schema for planetary exploration datasets, the Planetary Data System version 4 information model [^7] https://pds.nasa.gov/datastandards/ (PDS4-IM) (Hughes et al., 2017); and the OGC (Open Geospatial Consortium [^8] https://www.ogc.org/), which is commonly used for the study of planetary surfaces.
@@
The heliophysics community is proposing a set of tools and standards for finding and accessing datasets for the heliophysics community. Semantic artefacts in this community were historically of three kinds: the SPASE Ontology (available as an XML schema), which includes lists of terms, properties and classes for defining various objects and resources (Persons, Observatories, Instruments, Datasets, Repositories, etc); the VSO (Virtual Solar Observatory) metadata; and the SOLARNET set of keywords (dedicated to Solar observations).
@@
The IVOA, IPDA and IHDEA alliances are all worldwide working groups, consensus and bottom-up driven, and based on best effort contributions. Interdisciplinary links between these communities have been mostly developed thanks to the Europlanet/VESPA [^9] http://www.europlanet-vespa.eu infrastructure (Erard et al., 2018). They focus on discoverability and implementation of plugins to extend the capabilities of existing tools. The semantic interoperability across the sub-communities approach started only recently, with the ongoing development of two common SAs: a vocabulary for “observation facilities” and another one for “reference frames”.
@@
The main challenge of this project is to produce SAs in a standard semantic web form (typically OWL — Ontology for the Web Language (Bechhofer et al., 2004), or SKOS — Simple Knowledge Organisation System (Miles and Bechhofer, 2009)). Besides being a requirement for being ingested into the astronomy OntoPortal instance, it is also a logical evolution for our artefacts. Indeed, most of the current SAs (except those from the celestial astronomy community) are in diverse forms, from lists of terms embedded in XML schemas, to unformatted lists of metadata in specification documents (PDF files, HTML pages, etc.) This “semantic lifting” shall be done by the semantic working groups or authorities of the relevant communities (e.g., the IVOA Semantics WG or the IHDEA dedicated teams), with the support of expert in semantics and partners from the OntoPortal Alliance (mostly from the AgroPortal team, the Agri-food community OntoPortal instance developed by INRAE).

### v3

@@ Abstract
The astronomy communities are widely recognised as mature communities for their open science practices. However, while their data ecosystems are rather advanced and permit efficient data interoperability, there are still gaps between these ecosystems. Semantic artefacts (SAs) —e.g., ontologies, thesauri, vocabularies or metadata schemas— are a means to bridge that gap as they allow to semantically described the data and map the underlying concepts. The increasing use of SAs in astronomy presents challenges in description, selection, evaluation, trust, and mappings. The landscape remains fragmented, with SAs scattered across various registries in diverse formats and structures — not yet fully developed or encoded with rich semantic web standards like OWL or SKOS — and often with overlapping scopes. Enhancing data semantic interoperability requires common platforms to catalog, align, and facilitate the sharing of FAIR (Findable, Accessible, Interoperable and Reusable) SAs. In the frame of the FAIR-IMPACT project, we prototyped a SA catalogue for astronomy, heliophysics and planetary sciences. This exercise resulted in improved vocabulary and ontology management in the communities, and is now paving the way for better interdisciplinary data discovery and reuse. This article presents current practices in our discipline, reviews candidate SAs for such a catalogue, presents driving use cases and the perspective of a real production service for the astronomy community based on the OntoPortal technology, that will be called OntoPortal-Astro.
@@ 1 Introduction
The set of resources in the so-called “virtual observatory” (Genova et al., 2000; Hill et al., 2009; Hanisch et al., 2015; Erard et al., 2018, 2020; Mampaey et al., 2025) forms a network of related data products, catalogues, services and tools allowing researchers to discover, access and process data for their research. One example is the International Virtual Observatory Alliance [^1] https://ivoa.net (IVOA) framework (Arviset et al., 2018), which proposes a registry (Demleitner et al., 2019) of resources gathering endpoints for services, datasets and catalogues, and is used by tools to discover and provide access to resources. This astronomy framework is an example of an open science ecosystem.
1.1 Astronomy communities
@@
The astronomy community is mainly composed of three semantics sub-communities: celestial astronomy [^2] In this paper, we use the term “celestial astronomy” to refer to any topics outside the Solar System. Note that it’s not a commonly accepted definition for this concept, but we use it for the sake of simplicity. (objects are referenced to with their fixed sky coordinates, e.g., stars, galaxies, cosmology, etc); heliophysics (the study of the Sun, and the plasma environments throughout the Solar System); planetary sciences (the study of the Solar System objects, e.g., planets, comets, asteroids, etc). Each of these sub-communities have developed interoperability frameworks and semantic ecosystems, which are rather siloed up to now, compromising the semantic and technical interoperability of those ecosystems.
@@
The main astronomy sub-community is the celestial astronomy community. Its data management infrastructure is organised around the IVOA, which is interoperability driven. The IVOA was founded in 2002, and is now composed of 23 national groups (at the time of writing), which act as relays between the community developments and the international cooperation level. The IVOA gathers twice per year with the so-called “IVOA Interop” conferences. With this fully bottom-up framework, the IVOA defines schemas, protocols and vocabularies, which are respectively managed by the Data Models, Data Access Layers, and Semantics working groups (WGs). In addition, a registry (managed by the IVOA Registry WG) provides a catalogue of IVOA compliant resources. Hence, this enables a stable and well-established ecosystem adopted by the vast majority of astronomy archives, data centres and tool developers. As a result, scientists are using applications, such as Aladin (Bonnarel et al., 2000) or TOPCAT (Taylor, 2005, 2011), which are fully relying on IVOA tooling without seeing any detail of the underlying interoperability layers.
@@
The second sub-community is the heliophysics community. It is organised around the International Heliophysics Data Environment Alliance [^3] https://ihdea.net (IHDEA, Masson et al., 2024), which is distribution [^4] In the sense of DCAT (Data Catalog Vocabulary), see https://www.w3.org/TR/vocab-dcat-3/#Class:Distribution driven. The IHDEA was formed in 2018 by nine institutions (from laboratories to space agencies) willing to foster collaboration and standardisation of the heliophysics data ecosystem. However, the most of heliophysics community standards predates the inception of the IHDEA. The Space Physics Archive Search and Extract (SPASE, Roberts et al., 2018) is the main heliophysics information model, which development started in 2002. It is maintained by the SPASE Metadata Working Team. It defines a rather complete and evolutive way to describe space physics observational datasets, model runs, instruments or repositories. For each dataset, it is possible to define the various ways of the accessing the data. It is used by several major heliophysics data centres (NASA/SPDF, CNES/CDPP, IUGONET, see e.g., Abe et al., 2014; Clarke, 2016; Roberts et al., 2018) for their internal data model. Other metadata guidelines are available to this community (such as the ISTP metadata for data stored in CDF (Whipple and Lancaster, 1995; Candey, 1998), or the SOLARNET metadata for data stored in FITS (Hanisch et al., 2001; Haugan and Fredvik, 2021)). Data access protocols are also standardised, with, e.g., the HAPI protocol (Weigel et al., 2021). The IHDEA ecosystem is not as interoperable as the IVOA one. While some tools are providing a very simple access to heliophysics datasets (such as the AMDA [^5] https://amda.irap.omp.eu (Génot et al., 2021) or the CDAWeb [^6] https://cdaweb.gsfc.nasa.gov web interfaces), using IHDEA services or resources often require a deep understanding of interoperability interfaces and data formats. However ongoing developments are aiming at building tighter links with IVOA and improve open science in heliophysics.
@@
The third sub-community is the planetary science community. It is mainly organised around the International Planetary Data Alliance [^7] https://planetarydata.org (IPDA), which is archive driven. The IPDA was formed in 2006 and is now composed of 13 members, which are mostly space agencies. The primary role of the IPDA is to develop and maintain the NASA Planetary Data System (PDS) (Hughes et al., 2017), which is an advanced information model for planetary exploration science archives. The PDS is used by all major space agency archive with planetary exploration missions: NASA/PDS, ESA/PSA, JAXA, ISRO, etc. Moreover, planetary scientists studying the planetary surfaces are using the OGC888 https://www.ogc.org/ (Open Geospatial Consortium) ecosystem, developed by the geoscience community (Earth observation). In 2021, a small group of planetary scientist formed the Planetary Domain Working Group, which role is to adopt the OGC standards for the study of planetary surfaces. Technical interoperability between OGC and IVOA has been studied for some years (Hare et al., 2018; Marmo et al., 2018; Minin et al., 2019; Erard and the VESPA team, 2024). However, semantic interoperability still needs to be implemented.
Each astronomy sub-community data framework implements various pieces of an open science ecosystem with a different goal: celestial astronomy focuses on interoperability; heliophysics, on distribution; and planetary sciences, on long term archive. This also implies that different ways of registering, connecting, publishing and accessing resources and their metadata.
1.2 Linked-data architecture
The linked-data architecture has been defined at the inception times of the World Wide Web (Berners-Lee and Hendler, 2001). The original framework — called RDF: Resource Description Framework — (Schreiber and Raimond, 2014) developed and published by the W3C proposes a description of resources and their relations, based on triples: a subject, a predicate and an object. The triples can be seen as sentences, the predicate being the verb. The set of triples creates an oriented and tagged graph of nodes. Nodes are resources (referred to as URI: Unified Resource Identifier). Predicates (or properties) designate the links connecting the ressources. The classes, their instances and the properties are usually defined in standard vocabularies or ontologies capturing a certain level of semantics that a machine will then rely to query and reason over the data.
In the present case, this graph of nodes, metadata and relations is capturing knowledge about scientific resources, hence we call it a “scientific knowledge graph”. Keeping this graph of resources usable and interoperable while it grows, requires semantic and technical coordination and synchronisation, so that terms are globally understood by the tools and actors providing the content and enabling the access. This semantic and technical coordination is achieved by methods, technologies and tools coming from a domain called, the Semantic Web for which semantics artefacts, are at the core (Corcho et al., 2024). A semantic artefact [^9] In Corcho et al. (2024), a “semantic artefact” is defined as a machine-actionable formalisation (represented using appropriate formats and serialisations, including RDF and non-RDF standards) of a conceptualisation, enabling sharing and reuse by humans and machines. In this paper, the term is used to refer to both machine-actionable and non-machine-actionable resources. (hereafter referred to as SA) — a broader term to include ontologies, terminologies, taxonomies, thesauri, vocabularies, metadata schemas — can be a simple list of terms (properties, classes of resources), or a more complex ontology allowing inference and logics between terms. We know that a unique absolute ontology is utopia. Each SAs encode contextual knowledge, which need to be understandable, shareable and reusable by humans and machines.
Without fully implementing yet the standard semantic web toolkit, the astronomy communities are implementing open linked-data frameworks. They all define resource classifications, propose community registries that provide links between the resources, and associate these with rich metadata. Upgrading these open science frameworks with semantic web technologies is then a matter of defining and mapping concepts without breaking the current solutions.
1.3 Semantics Artefacts for FAIR Astronomy
@@
The scope of this study covers the astronomy community along with its three sub-domains (celestial astronomy, planetary sciences and heliophysics). Recent publications (e.g., Hurlburt and Timmons, 2022; Berriman, 2024; Cecconi, 2024; Landais et al., 2024; Servillat et al., 2024; Masson et al., 2024) show that each of these communities are claiming that they are producing a FAIR (Findable, Accessible, Interoperable and Reusable, Wilkinson et al., 2016) ecosystem. The evaluation of the FAIRness of astronomy digital resources is discussed and studied by many groups, and is out of the scope of this paper. However, one of the FAIR criteria [^10] I2: “(meta)data use vocabularies that follow FAIR principles” explicitly requires that community vocabularies also complies with the FAIR principles. Moreover, the fact that each sub-community is still siloed is a strong hint that things can be improved.
In 2018, the IVOA Solar System interest group started to bring heliophysics and planetary sciences needs in the IVOA ecosystem. The uptake of IVOA protocols in heliophysics and planetary sciences is growing, but at lot remains to be done, in particular on the semantic interoperability side, so that it is more widely adopted. This means that we need to harmonise our SAs, so that communities can exchange data and transform these data into knowledge — semantically described, interoperable, actionable, and open. Moreover, the adoption of common SAs across communities will improve the FAIRness of data products, since the metadata would be described with the same terms, or with terms that are related one to the other. The increasing use of SAs in astronomy presents challenges including their description, selection, evaluation, trustworthiness and mapping. The landscape remains fragmented, with SAs scattered across various registries in diverse formats and structures — not yet fully developed or encoded with rich semantic web standards based on RDF like OWL (Ontology for the Web Language, Bechhofer et al. (2004)) or SKOS (Simple Knowledge Organisation System, Miles and Bechhofer (2009)) — and often with overlapping scopes. There is thus need for a common place to catalogue, manage, assess, select and align FAIR SAs from several communities (Gray, 2015; Ritschel et al., 2016; Rovetto, 2023).
@@ 2 Current semantic artefact management in astronomy
The celestial astronomy community data alliance is the IVOA, which maintains an operational interoperability framework used by data repositories and science application platforms throughout the world. In this community, SAs are composed of vocabularies and schemas. The Semantics WG of the IVOA is managing the vocabularies used in the IVOA standards, following the Vocabularies in the VO specification (Demleitner et al., 2023). The vocabularies are listed in a dedicated web page [^11] https://ivoa.net/rdf/ where they are accessible using IVOA as well as with semantic web standard formats. Updates of the vocabularies are managed through a Vocabulary Enhancement Proposal (VEP) process (for details, see Demleitner et al., 2023). The maintenance of the vocabularies is managed in a dedicated Github repository [^12] https://github.com/ivoa-std/Vocabularies. The Data Model WG of the IVOA is managing the schemas, following the Virtual Observatory Data Model Language specification (Lemson et al., 2018), but they are not available in semantic web compliant formats. The role on SAs in the IVOA is to ensure the interoperability, with controlled vocabularies to be used in protocol and metadata fields. In the recent years, most of the controlled valued lists, which were previously maintained in schemas by the Data Model WG, have been transferred to the Semantics WG. In this way, the IVOA SAs have already been prepared for a wider use.
@@
In the planetary science community, two main frameworks co-exist with different scopes. The IPDA maintains an advanced data archiving metadata schema for planetary exploration datasets, the Planetary Data System version 4 information model131313 https://pds.nasa.gov/datastandards/ (PDS4-IM) (Hughes et al., 2017). The foundations of the PDS4-IM rely on the OAIS Standard (Open Archive Information System, ISO 14721:2012). It is developed using semantic web related tools, such as Protégé (Musen, 2015), but is not implemented using RDF. The PDS4-IM is maintained and curated by a Change Control Board [^14] https://pds-engineering.jpl.nasa.gov/pds4/ccb/, which is supported from a Data Design WG [^15] https://pds-engineering.jpl.nasa.gov/pds4/ddwg/ representing the various PDS4 disciplinary nodes. The Planetary Domain WG of the OGC (Open Geospatial Consortium) proposes a set of standards and tools used for the study of planetary surfaces. Although the OGC already implements SAs in the form of RDF SAs (see, e.g., the ESIP —Earth Science Information Partners— Community Ontology Repository [^16] http://cor.esipfed.org), the Planetary Domain WG is not providing its SAs in RDF yet.
@@
The heliophysics data ecosystem is managed by the IHDEA. This community is proposing a set of tools and standards for finding and accessing datasets for the heliophysics community. Semantic artefacts in this community were historically of three kinds: the SPASE Ontology (available as an XML schema), which includes lists of terms, properties and classes for defining various objects and resources (Persons, Observatories, Instruments, Datasets, Repositories, etc); the VSO (Virtual Solar Observatory) metadata; and the SOLARNET set of keywords (dedicated to Solar observations). Each of these three metadata standard are community effort to improve the findability and accessibility of their research products and their distributions. However, they are still currently not fully connected, and none of the them share an RDF implementation.
@@
The IVOA, IPDA and IHDEA alliances are all worldwide collaborative groups, consensus and bottom-up driven, and based on best effort contributions. Interdisciplinary links between these communities have been mostly developed thanks to the Europlanet/VESPA [^17] http://www.europlanet-vespa.eu infrastructure (Erard et al., 2018). They focus on discoverability and implementation of plugins to extend the capabilities of existing tools. The semantic interoperability across the sub-communities approach started only recently, with the ongoing development of three common SAs: a vocabulary for “observation facilities” (Cecconi et al., 2023) and a proposal to the IHDEA to adopt two IVOA SAs for managing reference frames (Weigel et al., 2025a) and units (Weigel et al., 2025b).
In this paper, we present the results of a prototyping exercise. Thanks to the FAIR-IMPACT project, we have implemented an instance of the OntoPortal technology (Jonquet et al., 2023) to deploy a prototype SA catalogue for astronomy. Hereafter, we present current practices in astronomy, and review candidate SAs for such a cataloguing exercise. We report the use cases that have been driving this development. We present the first results, discuss the impact and outcomes, and layout a roadmap for the deployment of a real production service for the astronomy community based on the OntoPortal technology, that will be called OntoPortal-Astro. Throughout the paper, the astronomy OntoPortal instance and the OntoPortal-Astro terms refer respectively to the prototype instance development within the FAIR-IMPACT project, and the upcoming official server.
@@
The two main challenges of this project are: (i) to identify the SAs, which are currently scattered in many places, and (ii) to produce SAs in a standard semantic web form (typically OWL, or SKOS). Besides being a requirement for being ingested into the astronomy OntoPortal instance, it is also a logical evolution for our artefacts. Indeed, most of the current SAs (except those from the celestial astronomy community) are in diverse forms, from lists of terms embedded in XML schemas, to unformatted lists of metadata in specification documents (PDF files, HTML pages, etc.) This “semantic lifting” shall be done by the semantics-related working groups or authorities of the relevant communities (e.g., the IVOA Semantics WG or the IHDEA dedicated teams), with the support of expert in semantics and partners from the OntoPortal Alliance (mostly from the AgroPortal team, the Agri-food community OntoPortal instance developed by INRAE) as well as from the OntoPortal-Astro team (see Section 6).